# Wilsonian Renormalization of Neural Network Gaussian Processes

### v2

@@ I Introduction
Deep neural networks (DNNs) are showing an increasing amount of universality across tasks [1], dataset sizes, and computational power [2]. In simple asymptotic limits, power-law scaling of DNN performance [2] can be traced to Zipf’s law behaviour on the underlying datasets [3, 4]. In other, more non-linear limits, the effects of non-linearities can often be absorbed as a change of a single scale parameter [5, 6, 7]. As DNNs are large-scale systems in which the importance of each specific parameter is minute [8], some self-averaging effects leading to universality are perhaps expected, yet their extent remains to be explored both theoretically and empirically.
@@
In physics, scale-free behavior, renormalization of parameters, and universality, are key aspects of the Renormalization Group (RG), one of the pillars of theoretical physics [9, 10]. At its core, RG is an analytical technique that generates a continuous transformation (flow) between models/physical systems. This transformation is defined such that all models along a given flow agree on low-resolution/coarse-grained observables. Critical theories, showing many of the aforementioned scale-free phenomena, are described by points on this flow at which the correlation length diverges. Besides providing an analytical toolkit for tackling large-scale interacting systems (e.g., [11, 12, 13]), RG is an inherent part of modelling systems (via notions of relevancy), and offers qualitative insights into the role of non-linearities via the regularities of low-dimensional flows.
@@
In the context of increasing efforts to make AI safer and more predictable, the aforementioned scale-free universal behaviors of DNNs are encouraging and invite interpretations in terms of some renormalization group flow. At a practical level, there are various different procedures for implementing RG, the choice of which is best determined by physical criteria, chiefly: (i) What is the scale? This determines which observables are being integrated or coarse-grained out, and which remain. (ii) What is the UV theory on which one is applying RG? (iii) Is the induced RG flow tractable, or does it make some quantities more tractable?
@@
An early line of work sought to augment (or interpret) the action of a layer in deep neural networks so that it coarse-grains the data distribution, thus drawing an analogy with RG [14, 15, 16]. Similarly, the action of a diffusion model can be interpreted as inverting an RG flow on the data distribution [17]. The UV theories here are thus the data distribution and the IR theory consists of the relevant features in the data for the layer action and white noise data for the diffusion action. Another line of work [18, 19] views the distribution on functions induced by a random DNN as the UV theory, and uses RG formulations to treat non-Gaussian (finite-width) effects [20, 21, 22] in this distribution. However, the notion of scale and the relation to actual DNN priors remains quite varied across these works. Finally, another set of works relating RG flows and Bayesian Inference view both as forms of augmenting a probability distribution [23, 24, 25, 26, 27]. In this approach however, reducing data generates the RG flow, so the true data-generated model (or target function) is the UV theory, while the IR theory is a random DNN, so that observables that are consistently reproduced along the flow are the most unlearnable modes. To the best of our understanding, these works have yet to analytically predict the generalization performance of neural networks.
@@
Here we formulate a distinct route where RG may come into quantitative use in predicting DNN performance. Our UV theory is the distribution over outputs generated by an ensemble of trained neural networks at finite amount of data, while the IR is the theory for the learnable modes. To generate an RG flow, we integrate out unlearnable output modes and track their renormalizing effect on the learnable modes. To illustrate our approach, we focus on a suitable infinite-overparameterization limit [28, 29, 30, 31, 32] (the Gaussian Process limit) which applies to a wide range of DNNs. We focus on using RG to treat the non-linearities induced by data-averaging/finite-sample-size effects rather than those induced by finite-width, though our formalism naturally extends to that case as well.
@@
We show that in various relevant scenarios, our RG procedure is tractable and results in a renormalized Gaussian Process (GP) theory. When the unlearnable modes are uncorrelated, in a manner to be made precise below, this amounts to a scalar renormalization of the loss or equivalently the ridge parameter, thus providing an RG interpretation for the results of [5]; this is developed in sec. III. When correlations between unlearnable modes are present, a further input-dependent/spatial weighting of the loss appears as well as potential shifts to the target; this is demonstrated in sec. IV. The resulting RG flow proceeds until it reaches the first learnable output modes. Thus, our lower RG scale (i.e., the infrared) naturally depends on the type and amount of data. In cases where unlearnable modes have sizable collective fluctuations, the RG flow ends in a weakly interacting regime amenable to perturbative methods [32]. Alternatively, it may end in a strongly interacting theory which, in generic high dimensional scenarios, is amenable to saddle point techniques [5]. Lastly, we point out that for power-law decaying kernel spectra, often found in real-world data, we recover scaling-law results [3], thereby establishing a concrete link between RG and scaling-laws.
@@
By relying on rather general formalism and introducing a down-to-earth practically motivated notion of RG, we hope to provide a stepping stone towards establishing notions of universality in deep learning.
@@ II Equivalent Kernel Limit of the Replica Action
To provide a relatively self-contained derivation, we first define Gaussian Process Regression, its relations to neural networks, and replica techniques used to obtain the average predictions over ensembles of datasets. In addition to being a well-established tool in machine learning, Gaussian Process Regression describes both the lazy learning or kernel regime of an infinite-width neural network [30, 29, 32], alongside extensions to finite networks via approaches detailed in [6, 33, 34].
The output of such random networks behaves as a Gaussian Process with zero mean and a covariance function $K(x,x^{\prime})$, called the kernel. Here, the training data, $x,x^{\prime}\in\mathbb{R}^{d}$, are drawn from a data measure [^1] By this we simply mean a measure over the data that admits an interpretation as a probability distribution. That is, let $\mathrm{d}\mu_{x}$ be a measure on $x\in\mathbb{R}^{d}$ such that for some probability distribution $p_{\mathrm{data}}(x)$, we have $\int\!f(x)\mathrm{d}\mu_{x}=\int\!f(x)p_{\mathrm{data}}(x)\mathrm{d}x$ for $f:\mathbb{R}^{d}\to\mathbb{R}$. In a slight abuse of notations, we will refer to both the measure $\mathrm{d}\mu_{x}$ and the distribution $p_{\mathrm{data}}(x)$ as the data measure. $p_{\mathrm{data}}(x)$; the total training set consists of $n$ such draws, and is denoted by the $n\times d$ matrix $X_{n}$. We further consider a regression problem with a target, $y(x)$. A standard result in GP regression states that [35] the average predictor on a test point, $\tilde{f}(x_{*})$, given $X_{n}$ is
@@
Since the average predictor (1) is conditional on $X_{n}\!\sim\!p_{\mathrm{data}}(x)$, it is sensitive to the details of the specific dataset we drew. To remove this dependence, we consider the dataset-averaged predictor by computing the expectation value with respect to $p_{\mathrm{data}}(x)$. The most general way of doing so in a tractable manner uses a “grand-canonical” trick [36, 32], which assumes that the number of training points is drawn from a Poisson distribution with expectation value $\eta$,
@@
This replica approach was first applied to GP regression in [36], and was used in [32] to obtain
@@
$S=$ $\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})$ (5)
$\qquad-\eta\int\!\mathrm{d}\mu_{x}\,e^{-\sum_{m=1}^{M}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2\sigma^{2}}}\leavevmode\nobreak\ ,$
@@
Although our initial field theory was Gaussian, averaging over the datasets resulted in a non-quadratic action, due to “quenched-disorder” induced by the training set. A simple yet useful limit [37] of the above theory, known as the Equivalent Kernel (EK) limit, is obtained when $\eta,\sigma^{2}\rightarrow\infty$ with $\eta/\sigma^{2}$ fixed. In this case the exponential on the right-hand side of (5) can be Taylor expanded, so that, at leading order, we obtain a non-interacting theory,
$\eta+S=$ $\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})$ (6)
$\qquad+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$
@@
see Appendix A. Note that the path integral measure also decouples into a measure over the feature modes, $\mathcal{D}f_{m}\sim\prod_{k}\mathcal{D}f_{mk}$, up to a constant Jacobian factor. Given this action, one can then show that the average prediction is [32, 35]
@@
$\mathrm{Var}[f_{k}]=\frac{1}{\lambda_{k}^{-1}+\eta\,\sigma^{-2}}=\frac{\sigma^{2}}{\eta}\frac{\lambda_{k}}{\lambda_{k}+\sigma^{2}/\eta}\leavevmode\nobreak\ .$ (12)
In the context of our RG prescription introduced below, these expression have the following important interpretation: modes with $\lambda_{k}\ll\sigma^{2}/\eta$ essentially behave as if $\sigma^{2}=\infty$ or $\eta=0$, i.e., they are decoupled from the inference problem and hence could just as well have been integrated out. This fact is already implicit in the action (10), where the term proportional to $\eta/\sigma^{2}$ becomes negligible compared to $1/\lambda_{k}$ for sufficiently small $\lambda_{k}$.
@@ III Renormalizing Gaussian features
We refer to the coefficients $f_{mk}$ as the GP modes, and the functions $\phi_{k}$ as the feature modes. Note that the former carry replica indices and may be highly correlated, while the feature modes are an inert property of the kernel/lazy-network. In the Equivalent Kernel limit, one sees that the action (10) splits cleanly into greater and lesser modes, $S=S_{0}[f_{m<}]+S_{0}[f_{m>}]$. At finite $\eta,\sigma^{2}$ however, the replicated action (5) will contain interaction terms collectively denoted $S_{\mathrm{int}}[f_{m<},f_{m>}]$. Explicitly, making the decomposition (13), (14) in the replicated action (5), we have
@@
$-\eta\int\!\mathrm{d}\mu_{x}\,\exp\!\Bigg{(}-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\left(f_{m<}-y_{<}\right)\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+\left(f_{m>}-y_{>}\right)\right]^{2}\Bigg{)}\leavevmode\nobreak\ ,$
where the first line and second line respectively define $S_{\mathrm{0}}[f_{m<}],\leavevmode\nobreak\ S_{\mathrm{0}}[f_{m>}],$ and $S_{\mathrm{int}}$. In general then, the exact partition function (4) takes the form
$\langle Z^{M}\rangle_{\eta}$ $=e^{-\eta}\int\!\prod_{m}\mathcal{D}f_{m<}\,e^{-S_{0}[f_{m<}]}$ (16)
×$\qquad\times\int\prod_{m}\mathcal{D}f_{m>}\,e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}$
@@
$S_{\mathrm{eff}}[f_{m<}]\coloneqq$ $-\ln\Bigg{(}e^{-S_{0}[f_{m<}]}\int\prod_{m}\mathcal{D}f_{m>}\,$ (17)
×$\qquad\qquad\times e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}\Bigg{)}\leavevmode\nobreak\ ,\leavevmode\nobreak$
@@
Our objective is now to perform the integral over the higher modes in (17). However, the presence of the interaction term complicates this, so we will first recast this term in a way which makes the interaction between the lesser and greater GP modes more tractable. To this end, we observe that the feature modes, $\phi_{k}(x)$, are functions over the data, which in turn is drawn from the data measure, $\mathrm{d}\mu_{x}=p_{\mathrm{data}}(x)\mathrm{d}x$. It will prove more convenient to instead treat these as fluctuating random variables over the data measure. To do this, we insert an identity of the given form into the integral over the data measure $\mathrm{d}\mu_{x}$ in the interaction term
@@
$S_{\mathrm{int}}$ $=-\eta\int\!\mathrm{d}\mu_{x}\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)$ (19)
×$\qquad\qquad\times\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]^{2}\right\}$
$=-\eta\int\!\mathcal{D}\varphi\,P[\varphi]\,$
×$\qquad\qquad\times\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\varphi_{k}\right]^{2}\right\}\leavevmode\nobreak\ ,$
@@
$P[\varphi]\leavevmode\nobreak\ \coloneqq$ $P[\varphi_{1},\varphi_{2},...]=\int\!\mathrm{d}\mu_{x}\,\prod^{\infty}_{k=1}\delta\left(\varphi_{k}-\phi_{k}(x)\right)\leavevmode\nobreak\ ,\qquad$ (20)
$\mathcal{D}\varphi\leavevmode\nobreak\ \coloneqq$ $\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\leavevmode\nobreak\ .$ (21)
@@
$=-\eta\int\!\mathcal{D}\varphi\,P[\varphi_{<},\varphi_{>}]\,\exp\!\Bigg{\{}-\frac{1}{2\sigma^{2}}\big{(}\Phi_{<}^{\top}\Phi_{<}+\Phi_{>}^{\top}\Phi_{>}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2\Phi_{<}^{\top}\Phi_{>}\big{)}\Bigg{\}}\leavevmode\nobreak$
$=-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}$
×$\times\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak\ ,$
@@
We can now proceed via a two-stage process, wherein we first integrate out the greater feature modes with the GP modes held fixed, and then integrate out the greater GP modes ($f_{mk}$ with $k>\kappa$). To this end, we approximate $P[\varphi]$ as a multivariate Gaussian distribution. We note in passing that a weaker requirement, namely $\Phi_{>}$ is Gaussian under the $p_{\mathrm{data}}(x)$ data measure, also suffices; cf. [5]. At a technical level, this assumption allows us to perform the spatial integration in $S_{\mathrm{int}}$ and recast the action in terms of $f_{mk}$ alone.
Recalling that the greater and lesser feature modes are by definition linearly uncorrelated under the data measure, the multivariate Gaussian approximation implies that they are also statistically independent. Consequently, $P[\varphi_{>}|\varphi_{<}]=P[\varphi_{>}]={\cal N}[0,\mathbb{1};\varphi_{>}]$. [^5] For brevity, here we assumed that $\int\mathrm{d}\mu_{x}\phi_{k>\kappa}(x)=0$, which holds in practice for many kernels. Due to the orthogonality of feature modes (cf. (8)), this also holds whenever the constant function is spanned by the lesser feature modes ($\phi_{k\leq\kappa}(x)$). The above integral over $\varphi_{>}$ now becomes Gaussian, and may be readily evaluated by expressing it in the form
@@
$=e^{-\frac{1}{2}\mathrm{tr}\ln\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)}$
×$\quad\times\exp\Bigg{\{}\frac{1}{2\sigma^{4}}\Phi_{<}^{\top}F\Big{(}\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\Big{)}^{-1}F^{\top}\Phi_{<}\Bigg{\}}\leavevmode\nobreak\ ,$

### v3

# thanks: Alphabetical ordering

@@ I Introduction
Deep neural networks (DNNs) are showing an increasing amount of universality across tasks bommasani2022opportunitiesrisksfoundationmodels , dataset sizes, and computational power kaplan2020scaling . In simple asymptotic limits, power-law scaling of DNN performance kaplan2020scaling can be traced to Zipf’s law behaviour on the underlying datasets bahri2021explaining ; bordelon2021spectrum . In other, more non-linear limits, the effects of non-linearities can often be absorbed as a change of a single scale parameter canatar ; LiSompolinsky2021 ; atanasov2024scaling . As DNNs are large-scale systems in which the importance of each specific parameter is minute molchanov2017pruning , some self-averaging effects leading to universality are perhaps expected, yet their extent requires both theoretical and empirical exploration.
@@
In physics, scale-free behavior, renormalization of parameters, and universality, are key aspects of the Renormalization Group (RG), one of the pillars of theoretical physics peskin1995 ; cardy1996scaling . At its core, RG is an analytical technique that generates a continuous transformation (flow) between models/physical systems. This transformation is defined such that all models along a given flow agree on low-resolution/coarse-grained observables. Critical theories, showing many of the aforementioned scale-free phenomena, are described by points on this flow at which the correlation length diverges. Besides providing an analytical toolkit for tackling large-scale interacting systems (e.g., WILSON1974 ; Gross1973 ; KT1973 ), RG is an inherent part of modelling systems (via notions of relevancy), and offers qualitative insights into the role of non-linearities via the regularities of low-dimensional flows.
@@
In the context of increasing efforts to make AI safer and more predictable, the aforementioned scale-free universal behaviors of DNNs are encouraging and invite interpretations in terms of some renormalization group flow. At a practical level, there are various procedures for implementing RG, the choice of which is best determined by physical criteria, chiefly: (i) What is the scale? This determines which observables are being integrated or coarse-grained out, and which remain. (ii) What is the UV theory on which one is applying RG? (iii) Is the induced RG flow tractable, or does it make some quantities more tractable?
@@
An early line of work sought to augment (or interpret) the action of a layer in deep neural networks so that it coarse-grains the data distribution, thus drawing an analogy with RG mehta2014exact ; Koch_Janusz_2018 ; Beny:2013pmv . Similarly, the action of a diffusion model can be interpreted as inverting an RG flow on the data distribution Cotler_2023 . The UV theories here are thus the data distribution and the IR theory consists of the relevant features in the data for the layer action and white noise data for the diffusion action. Another line of work Erbin_2022 ; erbin2023functional views the distribution on functions induced by a random DNN as the UV theory, and uses RG formulations to treat non-Gaussian (finite-width) effects Halverson:2020trp ; Grosvenor:2021eol ; Roberts_2022 in this distribution. However, the notion of scale and the relation to actual DNN priors remains quite varied across these works. Another set of works relating RG flows and Bayesian Inference view both as forms of augmenting a probability distribution berman2022dynamics ; berman2022inverse ; Berman_2023 ; berman2024ncoder ; howard2024bayesianrgflowneural . In this approach however, reducing data generates the RG flow, so the true data-generated model (or target function) is the UV theory, while the IR theory is a random DNN, so that the observables that are consistently reproduced along the flow are the most unlearnable modes. To the best of our understanding, these works have yet to analytically predict the generalization performance of neural networks. Finally, an earlier work Bradde_2017 models the data-distributions, irrespective of any neural network, as a field theory performs RG by removing large PCA components. Trading PCA with kernel-PCA, both that work and the current work share the same notion of UV and IR modes, however focus on very different ensembles/probability distributions.
@@
Here we formulate a distinct route where RG may come into quantitative use in predicting DNN performance. Our UV theory is the distribution over outputs generated by an ensemble of trained neural networks at finite amount of data, while the IR is the theory for the learnable modes. To generate an RG flow, we integrate out unlearnable output modes and track their renormalizing effect on the learnable modes. To illustrate our approach, we focus on a suitable infinite-overparameterization limit (the Gaussian Process limit) which applies to a wide range of DNNs, including state-of-the-art models such as transformers and Convolutional Neural Networks (CNNs) Neal ; jacot2018neural ; lee2018deep ; matthews2018gaussian ; Cohen_2021 . We focus on using RG to treat the non-linearities induced by data-averaging/finite-sample-size effects rather than those induced by finite-width, though our formalism naturally extends to that case as well.
@@
We show that in various relevant scenarios, our RG procedure is tractable and results in a renormalized Gaussian Process (GP) theory. When the unlearnable modes are uncorrelated, in a manner to be made precise below, this amounts to a scalar renormalization of the loss or equivalently the ridge parameter, thus providing an RG interpretation for the results of canatar ; this is developed in sec. III. When correlations between unlearnable modes are present, a further input-dependent/spatial weighting of the loss appears as well as potential shifts to the target; this is demonstrated in sec. V. The resulting RG flow proceeds until it reaches the first learnable output modes. Thus, our lower RG scale (i.e., the infrared) naturally depends on the type and amount of data. In cases where unlearnable modes have sizable collective fluctuations, the RG flow ends in a weakly interacting regime amenable to perturbative methods Cohen_2021 . Alternatively, it may end in a strongly interacting theory which, in generic high dimensional scenarios, is amenable to saddle point techniques canatar . Lastly, we point out that for power-law decaying kernel spectra, often found in real-world data, we recover scaling-law results bahri2021explaining , thereby establishing a concrete link between RG and scaling-laws.
@@
By relying on a rather general formalism and introducing a down-to-earth practically motivated notion of RG, we hope to provide a stepping stone towards establishing notions of universality in deep learning.
@@ II Equivalent Kernel Limit of the Replica Action
To provide a relatively self-contained derivation, we first define Gaussian Process Regression, its relations to neural networks, and replica techniques used to obtain the average predictions over ensembles of datasets. In addition to being a well-established tool in machine learning, Gaussian Process Regression describes both the lazy learning or kernel regime of an infinite-width neural network lee2018deep ; jacot2018neural ; Cohen_2021 , alongside extensions to finite networks via approaches detailed in LiSompolinsky2021 ; Seroussi2023 ; ariosto2022statistical .
The output of such random networks behaves as a Gaussian Process with zero mean and a covariance function $K(x,x^{\prime})$, called the kernel. Here, the training data, $x,x^{\prime}\in\mathbb{R}^{d}$, are drawn from a data measure [^2] By this we simply mean a measure over the data that admits an interpretation as a probability distribution. That is, let $\mathrm{d}\mu_{x}$ be a measure on $x\in\mathbb{R}^{d}$ such that for some probability distribution $p_{\mathrm{data}}(x)$, we have $\int\!f(x)\mathrm{d}\mu_{x}=\int\!f(x)p_{\mathrm{data}}(x)\mathrm{d}x$ for $f:\mathbb{R}^{d}\to\mathbb{R}$. In a slight abuse of notations, we will refer to both the measure $\mathrm{d}\mu_{x}$ and the distribution $p_{\mathrm{data}}(x)$ as the data measure. $p_{\mathrm{data}}(x)$; the total training set consists of $n$ such draws, and is denoted by the $n\times d$ matrix $X_{n}$. We further consider a regression problem with a target, $y(x)$. A standard result in GP regression states that Rasmussen the average predictor on a test point, $\tilde{f}(x_{*})$, given $X_{n}$ is
@@
Since the average predictor (1) is conditional on $X_{n}\!\sim\!p_{\mathrm{data}}(x)$, it is sensitive to the details of the specific dataset we drew. To remove this dependence, we consider the dataset-averaged predictor by computing the expectation value with respect to $p_{\mathrm{data}}(x)$. The most general way of doing so in a tractable manner uses a “grand-canonical” trick Malzahn ; Cohen_2021 , which assumes that the number of training points is drawn from a Poisson distribution with expectation value $\eta$,
@@
This replica approach was first applied to GP regression in Malzahn , and was used in Cohen_2021 to obtain
@@
$S=$ $\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})-\eta\int\!\mathrm{d}\mu_{x}\,e^{-\sum_{m=1}^{M}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2\sigma^{2}}}\leavevmode\nobreak\ ,$ (5)
@@
Although our initial field theory was Gaussian, averaging over the datasets resulted in a non-quadratic action, due to “quenched-disorder” induced by the training set. A simple yet useful limit Sollich2004 ; Silverman1984 of the above theory, known as the Equivalent Kernel (EK) limit, is obtained when $\eta,\sigma^{2}\rightarrow\infty$ with $\eta/\sigma^{2}$ fixed. In this case the exponential on the right-hand side of (5) can be Taylor expanded [^5] Note that while the Taylor expansion is exact in this limit, it is still quite a good approximation even at finite $\eta$ and $\sigma^{2}=O(0.01)$ Sollich2004 . It is also a well-known approximation Silverman1984 for the finite ridge case., so that, at leading order, we obtain a non-interacting theory,
$\eta+S=$ $\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$ (6)
@@
see Appendix A. Note that the path integral measure also decouples into a measure over the feature modes, $\mathcal{D}f_{m}\sim\prod_{k}\mathcal{D}f_{mk}$, up to a constant Jacobian factor. Given this action, one can then show that the average prediction is Cohen_2021 ; Rasmussen
@@
$\mathrm{Var}[f_{k}]=\frac{1}{\lambda_{k}^{-1}+\eta/\sigma^{2}}=\frac{\sigma^{2}}{\eta}\frac{\lambda_{k}}{\lambda_{k}+\sigma^{2}/\eta}\leavevmode\nobreak\ .$ (12)
In the context of our RG prescription introduced below, these expressions have the following important interpretation: modes with $\lambda_{k}\ll\sigma^{2}/\eta$ essentially behave as if $\sigma^{2}=\infty$ or $\eta=0$, i.e., they are decoupled from the inference problem and hence could just as well have been integrated out. This fact is already implicit in the action (10), where the term proportional to $\eta/\sigma^{2}$ becomes negligible compared to $1/\lambda_{k}$ for sufficiently small $\lambda_{k}$.
@@ III Renormalizing Gaussian features
In this section, we consider the case in which the distribution of feature modes, $P[\varphi]$ (to be made precise below), is Gaussian. This will help establish the notation which we will later use to generalize to the non-Gaussian case in section V. As we will later see, this Gaussian assumption is commonly made due in part to the fact that the feature distributions of many real-world datasets are (approximately) Gaussian. In section IV, we give a practical demonstration of this effect; namely, we show that our RG approach can predict the scaling of neural network performance with dataset size, matching both empirical observations and previous results obtained with alternative methods canatar ; bahri2021explaining .
@@
In the Equivalent Kernel limit, one sees that the action (10) splits cleanly into greater and lesser modes, $S=S_{0}[f_{m<}]+S_{0}[f_{m>}]$. At finite $\eta,\sigma^{2}$ however, the replicated action (5) will contain interaction terms collectively denoted $S_{\mathrm{int}}[f_{m<},f_{m>}]$. Explicitly, making the decomposition (13), (14) in the replicated action (5), we have
@@
$-\eta\int\!\mathrm{d}\mu_{x}\,\exp\!\Bigg{(}-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\left(f_{m<}-y_{<}\right)+\left(f_{m>}-y_{>}\right)\right]^{2}\Bigg{)}\leavevmode\nobreak\ ,$
where the first line defines the quadratic parts $S_{\mathrm{0}}[f_{m<}]$ and $S_{\mathrm{0}}[f_{m>}]$, and the second line defines the interaction part $S_{\mathrm{int}}$. In general then, the exact partition function (4) takes the form
$\langle Z^{M}\rangle_{\eta}$ $=e^{-\eta}\int\!\prod_{m}\mathcal{D}f_{m<}\,e^{-S_{0}[f_{m<}]}\int\prod_{m}\mathcal{D}f_{m>}\,e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}$ (16)
@@
$S_{\mathrm{eff}}[f_{m<}]\coloneqq-\ln\Bigg{(}e^{-S_{0}[f_{m<}]}\int\prod_{m}\mathcal{D}f_{m>}e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}\Bigg{)}\leavevmode\nobreak\ ,\leavevmode\nobreak$ (17)
@@
Our objective is now to perform the integral over the higher modes in (17). However, the presence of the interaction term complicates this, so we will first recast this term in a way which makes the interaction between the lesser and greater GP modes more tractable. To this end, we observe that the feature modes, $\phi_{k}(x)$, are functions over the data, which in turn is drawn from the data measure, $\mathrm{d}\mu_{x}=p_{\mathrm{data}}(x)\mathrm{d}x$. It will prove more convenient to instead treat these as fluctuating random variables over the data measure. To do this, we insert the identity in the following form into the integral over the data measure $\mathrm{d}\mu_{x}$ in the interaction term:
@@
$S_{\mathrm{int}}$ $=-\eta\int\!\mathrm{d}\mu_{x}\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]^{2}\right\}$ (19)
$=-\eta\int\!\mathcal{D}\varphi\,P[\varphi]\,\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\varphi_{k}\right]^{2}\right\}\leavevmode\nobreak\ ,$
@@
$P[\varphi]\coloneqq$ $P[\varphi_{1},\varphi_{2},...]=\int\!\mathrm{d}\mu_{x}\,\prod^{\infty}_{k=1}\delta\left(\varphi_{k}-\phi_{k}(x)\right)\leavevmode\nobreak\ ,\qquad$ (20)
$\mathcal{D}\varphi\coloneqq$ $\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\leavevmode\nobreak\ .$
@@
$=-\eta\int\!\mathcal{D}\varphi\,P[\varphi_{<},\varphi_{>}]\,\exp\!\Bigg{\{}-\frac{1}{2\sigma^{2}}\big{(}\Phi_{<}^{\top}\Phi_{<}+\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\big{)}\Bigg{\}}\leavevmode\nobreak$
$=-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak\ ,$
@@
We can now proceed via a two-stage process, wherein we first integrate out the greater feature modes with the GP modes held fixed, and then integrate out the greater GP modes ($f_{mk}$ with $k>\kappa$). To this end, we approximate $P[\varphi]$ as a multivariate Gaussian distribution. We note in passing that a weaker requirement, namely $\Phi_{>}$ is Gaussian under the $p_{\mathrm{data}}(x)$ data measure, also suffices; cf. canatar . At a technical level, this assumption allows us to perform the spatial integration in $S_{\mathrm{int}}$ and recast the action in terms of $f_{mk}$ alone.
Recalling that the greater and lesser feature modes are by definition linearly uncorrelated under the data measure, the multivariate Gaussian approximation implies that they are also statistically independent. Consequently, $P[\varphi_{>}|\varphi_{<}]=P[\varphi_{>}]={\cal N}[0,\mathbb{1};\varphi_{>}]$ [^7] For brevity, here we assumed that $\int\mathrm{d}\mu_{x}\phi_{k>\kappa}(x)=0$, which holds in practice for many kernels. Due to the orthogonality of feature modes (cf. (8)), this also holds whenever the constant function is spanned by the lesser feature modes ($\phi_{k\leq\kappa}(x)$).. The above integral over $\varphi_{>}$ now becomes Gaussian, and may be readily evaluated by expressing it in the form
@@
$=e^{-\frac{1}{2}\mathrm{tr}\ln\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)}\exp\Bigg{\{}\frac{1}{2\sigma^{4}}\Phi_{<}^{\top}F\Big{(}\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\Big{)}^{-1}F^{\top}\Phi_{<}\Bigg{\}}\leavevmode\nobreak\ ,$